A toy ball in the shape (Sphere) expands and contracts. When it is completely closed it has a diameter of (4.5. inches). Find the (Volume) of the (Sphere) when it is completely closed. (use 3.14 for pi)
Hint: the Volume of a sphere of radius r is 4/3 mr3)
The (Volume) of the (sphere) is approximately ?

Answer by CubeyThePenguin(3113)

V = (4/3) * pi * r^3

V = (4/3) * pi * (4.5/2)^3 = (4/3) * 3.14 * (2.25)^3 = 47.68 in^3